OpenStudy (anonymous):

how to prove c^2=a^2+b^2?

OpenStudy (amorfide):

let's say we have a triangle each side is length 3, 4 and 5cm 5 cm is the longest side so lets draw this triangle 

OpenStudy (amorfide):

C= longest side A= one of the other sides B= one of the other sides okay so C²=A²+B² 5²=3²+4² 25=9+16 25=25

OpenStudy (amorfide):

i think this would be enough to prove it... someone correct me if i am wrong

OpenStudy (anonymous):